# references.bib

### v4

@@ 4.5 Definition for transversally H-elliptic operators
Using Theorem 1.6.13, we see that the operator $K$ itself is compact. To prove that $(1-P^{2})e$ is compact for all $e\in C_{c}^{\infty}(G)$, let $F=d_{G}^{\ast}Kd_{G}$. As $\sigma_{H}(d_{G}^{\ast}d_{G})=\sigma_{H}(\Delta_{G})$, we prove that $\sigma_{H}(1-P^{2})-\sigma_{H}(F)$ is of negative order, such that $1-P^{2}$ is equal to $F$ modulo operators of negative order.
By Lemma 6.6 in [kasparov2017elliptic], for any arbitrary $e\in C_{c}(G)$, the representation of $e$ on $F$, $Fe=d_{G}^{\ast}Kd_{G}\circ e$ is equal to the product of the compact operator $d_{G}^{\ast}K$ and the bounded convolution operator with $d(e)$. Therefore, $Fe$ and thus $(1-P^{2})e$ is compact for arbitrary $e\in C_{c}(G)$, and by natural extension, for arbitrary $e\in C^{\ast}(G)$, establishing it as a $KK$-class in $K^{0}(C^{\ast}(G))$. ∎
@@
In general case that $M$ is locally compact, we ought to consider the index class in $K^{0}(C^{\ast}(G,C_{0}(M)))$ and symbol class in $\mathcal{R}KK(C_{0}(M),Cl_{\tilde{\Gamma}}(M)\otimes_{C(M)}C^{\ast}(Gr(H)))$ which is defined in [kasparov1988equivariant]. In the case that $M$ is compact, it can be simplified to our result by the induced map of the inclusion $\mathbb{C}\subset C(M)$ and the index class in $K^{0}(C^{\ast}(G))$ is called the distribution index of $P$.
@@
5.1 From elliptic to H-elliptic cases
In this section, we introduce the idea to prove the H-elliptic index theorem (Theorem 3.2.6), from the index theorem for elliptic operators.
Recall that the index theorem for elliptic operators (Theorem 2.1.4) is proved by using a rotation homotopy, which transforms multiplication variable $\xi$ in symbol $[\sigma(P)(x,\xi)]$ to the differential variable $\chi$ in the operator $P$. The Dolbeault element will be transformed into a $1$ element in $KK$-class. Such a homotopy induces a $KK$-product relation between $[P]$ and $[\sigma(P)]$ and leads to the index theorem.
Motivation for the proof
For an H-elliptic operator $P$ with conditions in theorem, in order to find such a relation represented by $KK$-product between the index class $[P]$ and symbol class $[\sigma_{H}(P)]$, it is most possible to find some similar homotopy to relate the representation elements of these two terms. However, there is a gap that $P$ is an operator on $M$ and $\sigma_{H}(P)$ is an operator on $Gr(H)$. The calculation for the H-cosymbol cannot be conducted on $TM$ or $T^{\ast}M$, preventing a direct computation of the H-cosymbol for both the symbol operator and the original operator in the same space, and thereby complicating the direct construction of a homotopy of rotation from the $\xi$-variable to $\chi$-variable. One possible way to do is to lift the operator $P$ to $Gr(H)$ and construct some homotopy on $Gr(H)$ to relate $P$ and $\sigma_{H}(P)$. Another possible way to analyze directly we found, is trying to analyze operators on the rectangle diagram
${Gr(H)}$${gr(H)}$${M}$${TM,}$
with the operator on $TM$ is the lift of $P$ to $TM$ and the operator on $Gr(H)$ is $\sigma_{H}(P)$, and embed them into this diagram. In this way, we could find some relation between the ordinary symbol $\sigma(P)$ and $\sigma(\sigma_{H}(P))$. However, $Gr(H)$ has no general coordinates, so we don’t know how to construct a natural homotopy on $Gr(H)$ or in such a diagram, and it’s difficult to deal with the Connes-Thom class directly because it can’t be represented explicitly. So we changed these approaches to the following, using the idea in the proof of the H-elliptic index theorem by Mohsen in [mohsen2022index]:
Main idea
We transform the symbol class from $K_{0}(C^{\ast}(Gr(H)))$ to $K^{0}(T^{\ast}M)$ by the composition of isomorphisms, see the composition $\mu_{s=0}\circ\mu_{t=0}^{-1}$ in the following diagram:
Note that it represent the $KK$-class $[\mathcal{E}]$, and $\mu_{t=0}$ is the Connes-Thom isomorphism. By Lemma 3.2.7, The $KK$-product of the symbol class $[\sigma_{H}(P)]$ and $[\mathcal{E}]$ can be represented by the Fourier transform of a family of operators $P^{\prime}$ on $T_{x}M$ with $x\in M$. In this way, we delete the Connes-Thom term and obtain a multiplication operator $\sigma^{\prime}(P)$ with coordinates $(x,\xi)$ on $TM$. Then we can do a similar way in the proof of Theorem 9.1 in [kasparov2022k] to show the existence of homotopy relating $\sigma^{\prime}(P)$ and the lifting of $P$ to $TM$.
We will use this approach to give a proof of Theorem 3.2.6 in detail. It will subsequently facilitate the derivation of the main theorem through a similar process.
Proof of Theorem 3.2.6
Proof.
Assume that $P_{0}$ is an H-elliptic differential operator and $P$ is its bounded transformation. The proof is separated into the following five parts:
  • (i)
    First, we come back to $C(M)$ in place of $\mathbb{C}$. We have the symbol class $\sigma_{H}(P)\in KK_{0}^{G}(C(M),C^{\ast}(Gr(H)))$ and the index class $[P]\in K^{0}(C^{\ast}(G,C(M)))$.
    Utilizing the Dolbeault element $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$ and the Bott–Dirac element $[\mathcal{C}]\in KK_{G}(C(M),C(M))$, we want to prove with $KK$-theoretic argument that $[\mathcal{C}]\otimes_{C(M)}[P]$ and $[\sigma_{H}(P)]$ $\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]\otimes_{C_{0}(TM)}[\mathcal{D}_{M}]$ yield identical elements, where $[\mathcal{E}]\in KK_{G}(C^{\ast}(Gr(H)),$ $C_{0}(TM))$ represents the Connes-Thom class we defined in section 3.2.
    The orthonormal coordinates of $T(TM)$ are regarded to be $(x,\xi,\chi,\zeta)$ with $(x,\xi)$, the coordinate of the manifold $TM$, $\chi$ the differential coordinate of $x$, and $\zeta$ the differential coordinate of $\xi$.
  • (ii)
    By the basic construction of $KK$-products, the $KK$-product element $[\mathcal{C}]\otimes_{C(M)}[P]$ can be represented by $(\mathcal{H},Q_{1}=N_{1}(\mathcal{C}\hat{\otimes}1+N_{2}1\hat{\otimes}\widetilde{P}))$, where =$\mathcal{H}=\{L^{2}(\Lambda^{\ast}(T_{x}M))\}_{x\in M}\otimes_{C(M)}L^{2}(E)$, $\widetilde{P}$ denotes a $KK$-theoretic connection of $P$. Note that $\mathcal{C}$ represents the Bott-Dirac operator on $L^{2}(\Lambda^{\ast}(T_{x}M))$, the operator $\widetilde{P}$ can be taken as the lift of $P$ to $TM$ and $1$ represents the identity operator on $L^{2}(\Lambda^{\ast}(M))$. The operator $Q_{1}$ denotes a pseudo-differential operator on $TM$.
    By Lemma 3.2.7, the $KK$-product $[\sigma_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]$ is represented by freezing coefficients of $\sigma_{H}(P)$ to the unit element $e\in Gr(H)_{x}$ for each $x\in M$ and transforming the family of the operators we acquired on $gr(H)_{x}$ to a family of operators on $T_{x}M$, and then taking the Fourier transform on such family. This process leads to the operator $\sigma^{\prime}(P)(x,\xi)$ on $C_{0}(TM)$, which is the Fourier transform of the freezing coefficient of the H-principal part of $P$ to each point $x\in M$. Consequently, the $KK$-product element $[\sigma_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]\otimes_{C_{0}(TM)}[\mathcal{D}_{M}]$ is represented as $(\mathcal{H},Q_{2}=N_{1}^{\prime}(\widetilde{\mathcal{D}_{M}}+N_{2}^{\prime}1\hat{\otimes}\sigma^{\prime}(P)))$. $1\hat{\otimes}\sigma^{\prime}(P)$ is considered over $TM$ with coefficient $(x,\xi)$, $\widetilde{\mathcal{D}_{M}}$ denotes the Dolbeault operator on $\{L^{2}(\Lambda^{\ast}(T_{x}M))\}_{x\in M}\otimes_{C(M)}L^{2}(E)$ lifted to $\mathcal{H}$, and the second $KK$-product is taken over $C_{0}(TM)$. Thus, the product operator $Q_{2}$ is a pseudo-differential operator on $TM$.
  • (iii)
    Now we construct the rotation homotopy and show that these two product operators are homotopic through this homotopy. It suffices to prove the existence of a uniform homotopy between operators $\mathcal{C}$ and $\mathcal{D}_{M}$, $\widetilde{P}$ and $\sigma^{\prime}(P)$.
    Note that $\widetilde{\mathcal{D}_{M}}$ and $\mathcal{C}\hat{\otimes}1$ are operators on $TM$, the symbol of Dolbeault operator is
    $\sigma_{\mathcal{D}_{M}}(x,\xi,\chi,\zeta)=\text{ext}(-\chi+i\zeta)+\text{int}(-\chi+i\zeta)$
    and the symbol of $\mathcal{C}$ is
    $\sigma_{\mathcal{C}}(x,\xi,\chi,\zeta)=\text{ext}(\xi+i\zeta)+\text{int}(\xi+i\zeta).$
    So the rotation homotopy in the $(\xi,\chi)$ variable:
    $\xi\mapsto\cos t\ \xi-\sin t\ \chi;\ \chi\mapsto\sin t\ \xi+\cos t\ \chi,\ t\in[0,\pi/2].$
    transforms $\mathcal{C}\hat{\otimes}1$ to $\widetilde{\mathcal{D}_{M}}$.
  • (iv)
    By the argument in Remark 1.4.6, the $\sigma^{\prime}(P)$ is given by the Fourier transform of the freezing coefficient operators of the H-principal part of $P$ to each $x\in M$. The rotation homotopy above, which transforms the multiplication with $\xi$-variable by differential with $x$ variable in the operator sense, transforms the multiplication operator $\sigma^{\prime}(P_{0})(x,\xi)$, which is a polynomial in $\xi$, to an operator with the same H-principal part with the differential operator $\tilde{P_{0}}$, and thus, by functional calculus, transforms $\sigma^{\prime}(P)(x,\xi)$ to an operator having the same $KK$-representation with $\tilde{P}$ on $TM$. As a result, up to lower order terms, $\sigma^{\prime}(P)$ is homotopic to $\tilde{P}$ by the rotation homotopy.
    During the homotopy we can choose appropriate $N_{1}^{t},N_{2}^{t}$ in the $KK$-product construction, such that it fits with the rotation homotopy and satisfies $N_{1}^{0}=N_{1},N_{1}^{\pi/2}=N_{1}^{\prime},N_{2}^{0}=N_{2},N_{2}^{\pi/2}=N_{2}^{\prime}$.
    Based on the above argument, we observed that there is a rotation homotopy between the operators representing $[\mathcal{C}]\otimes_{C(M)}[P]$ and $[\sigma_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]\otimes_{C_{0}(TM)}[\mathcal{D}_{M}]$ in $K^{0}_{G}(C(M))$. As a result, they represent the same $KK$-class.
  • (v)
    Following Remark 10.4 in [kasparov2022k], the restriction map from $KK_{G}^{0}(C(M))$ to $KK_{G}^{0}(\mathbb{C})$, induced by the inclusion $\mathbb{C}\subset C(M)$, assigns the index class $[P]$ to the analytical index of $P$. As $[\mathcal{C}]$ represents the trivial element $1\in KK_{G}(C(M),C(M))$, the result follows if we take the restriction of these two $KK$-classes to $KK^{0}_{G}(\mathbb{C})$ induced by the inclusion map.
That proves Theorem 3.2.6.
5.2 Main index theorem for transversally H-elliptic differential operators
Recall Definition 2.2.7 of the Clifford–Dolbeault element $[\mathcal{D}^{cl}_{M,{\Gamma}}]\in K^{0}(C^{\ast}(G,$ $Cl_{{\Gamma}}(TM)))$. The $KK$-element $[\mathcal{D}^{cl}_{M,\tilde{\Gamma}}]$ is defined similar to $[\mathcal{D}^{cl}_{M,{\Gamma}}]$.
@@ 5.2 Main index theorem for transversally H-elliptic differential operators
The Clifford Dolbeault element $[D^{cl}_{M,\tilde{\Gamma}}]\in K^{0}(C^{\ast}(G,Cl_{\tilde{\Gamma}}(TM)))$ is defined to be the $KK$-element $[(\mathcal{H},\pi,F)]$, where
  • $\mathcal{H}=L^{2}(\Lambda_{\mathbb{C}}^{\ast}(TM))$;
  • $F=\mathcal{D}_{M}$ is the same operator as the Dolbeault element $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$;
  • The representation $\pi$ of $C^{\ast}(G,Cl_{\tilde{\Gamma}}(TM))$ is induced from the action $\zeta\mapsto\text{ext}(\psi^{-1}(\zeta))+\text{int}(\psi^{-1}(\zeta))$ for $\zeta$ a section of $\tilde{\Gamma}$, where $\psi$ denotes the bundle isomorphism defined in Definition 4.1.1.
For a transversally H-elliptic operator $P$, we obtain the symbol class
$[\sigma^{tcl}_{H}(P)]\in K_{0}^{G}(C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M)),$
@@
Let $M$ be a closed filtered Riemannian manifold and $G$ be a Lie group isometrically acting on $M$ and preserving the filtered structure. Let $P_{0}$ be a $G$-invariant differential operator on $M$ with its bounded transform
$P=\begin{pmatrix}0&P_{0}^{\ast}/(1+P_{0}P_{0}^{\ast})\\ P_{0}/(1+P_{0}^{\ast}P_{0})&0\end{pmatrix}$
being transversally H-elliptic of order 00, then we have
$[P]=j^{G}([\sigma^{tcl}_{H}(P)])\otimes_{C^{\ast}(G,C^{\ast}(Gr(H)))}j^{G}([\mathcal{E}])\otimes_{C^{\ast}(G,C_{0}(T^{\ast}M))}[\mathcal{D}_{M,\tilde{\Gamma}}^{cl}],$
where the $KK$-product on the right-hand side is considered
$KK(C^{\ast}(G),C^{\ast}(G,C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M))$
$\otimes_{C^{\ast}(G,C^{\ast}(Gr(H))))}$ $KK(C^{\ast}(G,C^{\ast}(Gr(H))),C^{\ast}(G,C_{0}(T^{\ast}M)))$
$\otimes_{C^{\ast}(G,C_{0}(T^{\ast}M))}$ $KK(C^{\ast}(G,Cl_{\tilde{\Gamma}}(M)\otimes C_{0}(T^{\ast}M)),\mathbb{C}).$
The proof of the H-elliptic index theorem, Theorem 3.2.6 in Section 3.2, employed the approach in [kasparov2022k] by $KK$-theory. Subsequently, the proof for transversally H-elliptic cases parallels the methodology used in the proof of Theorem 10.2 in [kasparov2022k] for transversally elliptic operators by Kasparov.

### v5

@@ 4.5 Definition for transversally H-elliptic operators
Using Remark 1.6.14, we see that the operator $K$ itself is of negative order. To prove that $a(1-P^{2})e$ is compact for all $a\in C_{c}^{\infty}(M)$ and $e\in C_{c}^{\infty}(G)$, let $F=d_{G}^{\ast}Kd_{G}$. As $\sigma_{H}(d_{G}^{\ast}d_{G})=\sigma_{H}(\Delta_{G})$, we see that $\sigma_{H}(1-P^{2})-\sigma_{H}(F)$ is of negative order, and thus $1-P^{2}$ is equal to $F$ modulo operators of negative order.
By Lemma 6.6 in [kasparov2017elliptic], for any arbitrary $e\in C_{c}(G)$, the representation of $e$ on $F$, $aFe=a\cdot d_{G}^{\ast}Kd_{G}\circ e$ is equal to the product of the compact operator $a\cdot d_{G}^{\ast}K$ and the bounded convolution operator with $d(e)$. Therefore, $aFe$ and thus $a(1-P^{2})e$ is compact for arbitrary $e\in C_{c}(G)$, and by natural extension, for arbitrary $a\in C_{0}(M)$ and $e\in C^{\ast}(G)$, establishing it as a $KK$-class in $K^{0}(C^{\ast}(G,C_{0}(M)))$. ∎
@@
In the case that $M$ is compact, the index element $[P]\in K^{0}(C^{\ast}(G,C_{0}(M)))$ can be simplified by the induced map of the inclusion $\mathbb{C}\subset C_{0}(M)$ and the resulting index class in $K^{0}(C^{\ast}(G))$ is called the distribution index of $P$.
@@
In this chapter, we give the main index theorem for transversally H-elliptic operators and prove it using a similar approach in [kasparov2022k] and [kasparov2024coarsepseudodifferentialcalculusindex].
5.1 Main index theorem for transversally H-elliptic operators
We denote the algebra $Cl_{\tau}(M)\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M)$ by $Cl_{\tau\oplus\Gamma}(M)$. First, we define the Dirac element with Clifford action and use it to define the Clifford-Dolbeault element similar to the construction we have done in Definition 3.3.4.
@@ 5.1 Main index theorem for transversally H-elliptic operators
The Dirac element $[d_{M,\Gamma}]\in K^{0}(C^{\ast}(G,Cl_{\tau\oplus\Gamma}(M)))$ is given by the pair $(H,F_{M})$, where $H=L^{2}(\Lambda^{\ast}(M))$, the action of $Cl_{\tau\oplus\Gamma}(M)$ on $H$ is defined on (real) covectors by
$\xi_{1}\oplus\xi_{2}\mapsto\text{ext}(\xi_{1})+\text{int}(\xi_{1})+i(\text{ext}(\xi_{2})-\text{int}(\xi_{2})).$
Here $\xi_{1}$ is a section of $\tau=TM$ and $\xi_{2}$ is a section of $\Gamma$. The representation of $C^{\ast}(G,Cl_{\tau\oplus\Gamma}(M))$ on $H$ is induced by the covariant representation defined by the action of $G$ and the above action of $Cl_{\tau\oplus\Gamma}(M)$. We take $D_{M}=d_{M}+d^{\ast}_{M}$ and let $F_{M}:=D_{M}(1+D_{M}^{2})^{-1/2}$.
Definition 5.1.2.
The Clifford-Dolbeault element $[{\mathcal{D}_{M,\Gamma}^{cl^{\prime}}}]\in K^{0}_{G}(C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}$ $Cl_{\Gamma}(M)))$ is the $KK$-product of the element $j^{G}([\eta_{C^{\ast}(Gr(H))}]\otimes_{C_{0}(M)}1_{Cl_{\Gamma}(M)})\in\mathcal{R}KK(M;$ $C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M),Cl_{\tau\oplus\Gamma}(M))$ and the Dirac element $[d_{M,\Gamma}]\in K_{G}^{0}(C^{\ast}(G,Cl_{\tau\oplus\Gamma}(M)))$.
Remark 5.1.3.
This Clifford-Dolbeault element is different from the one we defined in Definition 2.2.7. In fact, similar to the argument in Remark 3.3.5, $[\mathcal{D}_{M,\Gamma}^{cl}]$ is actually the $KK$-product of the fiberwise Dirac and the Dirac element with respect to the Clifford action (see Definition 8.17 in [kasparov2017elliptic]), and the new Clifford-Dolbeault element $[{\mathcal{D}_{M,\Gamma}^{cl^{\prime}}}]$ is defined replacing the fiberwise Dirac by the global dual Dirac element $\eta$.
For a transversally H-elliptic operator $P$, by Definition 4.5.6 and Theorem 4.5.7, we have the tangent-Clifford symbol class
$[\sigma^{tcl}_{H}(P)]\in KK^{G}(C_{0}(M),C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{{\Gamma}}(M)),$
@@
Let $M$ be a complete filtered Riemannian manifold and $G$ be a locally compact Lie group properly and isometrically acting on $M$ and preserving the filtered structure. Let $P$ be a $G$-invariant transversally H-elliptic operator on $M$ of order 00, then we have
$[P]=j^{G}([\sigma^{tcl}_{H}(P)])\otimes_{C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M))}[\mathcal{D}_{M,{\Gamma}}^{cl^{\prime}}],$
where the $KK$-product on the right-hand side is considered over
$KK(C^{\ast}(G,C_{0}(M)),C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{{\Gamma}}(M))$
$\otimes_{C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M))}KK(C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M)),\mathbb{C}).$
5.2 Some preliminaries for the proof
5.2.1 Local homeomorphisms
Remember that there is an exponential map $\exp_{x}$ which maps a small open ball $V_{x}\subset\tau_{x}=T_{x}M$ with center $0\in\tau_{x}$ diffeomorphically onto a small neighborhood $U_{x}$ of the point $x\in M$. The tangent map $(\exp_{x})_{\ast}$ is close to an isometry when $V_{x}$ is sufficiently small, and the smaller $V_{x}$ is, the closer to an isometry this map is. For any $x\in M$, there is also a similar exponential map $\text{Exp}_{x}:gr(H)_{x}\to Gr(H)_{x}$ which is a diffeomorphism of a small neighborhood of $0\in gr(H)_{x}$ to a small neighborhood of $0\in Gr(H)_{x}$. By the identification of $gr(H)_{x}$ with $\tau_{x}$ by the isomorphism, we can choose the neighborhood $V_{x}$ small enough so that $\text{Exp}_{x}$ maps this neighborhood $V_{x}\subset gr(H)_{x}$ to some neighborhood $\mathcal{V}_{x}\subset Gr(H)_{x}$. We can identify all three $L^{2}$-spaces: $L^{2}(U_{x})$, $L^{2}(V_{x})$ and $L^{2}(\mathcal{V}_{x})$ by the isometric isomorphisms.
5.2.2 The construction from $\sigma_{H}(P)$ to $P$
Let $\sigma_{H}$ be a cosymbol in $\mathcal{L}(E)\otimes\mathcal{M}(C^{\ast}(Gr(H)))$. Choose a continuous family of open balls $\{U_{x},x\in M\}$ which satisfy the following properties: the vector bundle $E$ is trivial over each $U_{x}$ (and this trivialization depends continuously on $x$), and the maps $\exp^{-1}_{x}:U_{x}\to V_{x}$ and $\text{Exp}_{x}:V_{x}\to\mathcal{V}_{x}$ are well defined diffeomorphisms.
Pick a function $\nu\in C_{0}([0,1))$ such that $0\leq\nu\leq 1$, $\nu(t)=1$ for $t\leq 1/2$, and $\nu(t)=0$ for $t\geq 2/3$. Assume that $V_{x}\subset\tau_{x}$ is a Euclidean ball of radius $r_{x}$, define the function $\nu_{x}(v)$ on $V_{x}$ by $\nu_{x}(v)=\nu(\lVert x\rVert/r_{x})$.
Using the isomorphisms $L^{2}(U_{x})\cong L^{2}(V_{x})\cong L^{2}(\mathcal{V}_{x})$, we can transplant $\nu_{x}$ into $L^{2}(\mathcal{V}_{x})$, then transplant the operator $\nu_{x}\sigma_{H,x}\nu_{x}$ into $\mathcal{L}(E_{x}\otimes L^{2}(U_{x}))$. This operator will be called $P(x)$.
[…]